# On the third homology of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$

### v1

@@ 2. Low dimensional homology of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$
where $\mu_{2}(A)=\{a\in{A^{\times}}:a^{2}=1\}$ and $i:\mu_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A)$ is given by $a\mapsto aI_{2}={\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a\end{array}\!\!\right)}$.
@@
Let $A$ be a ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu_{2}(A)\simeq\mu_{2}(F)$. Then we have the exact sequence
@@
First we study the homomophism $\tau:\Gamma(\mu_{2}(A))\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$. This map is induced by the quadratic map
@@
where $i_{\ast}:\mu_{2}(A)\simeq H_{1}(\mu_{2}(A),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is induced by $i$. Since $w^{2}=-I_{2}$, we have
$(-1)\otimes i_{\ast}(-1)=(-1)\otimes\overline{[-I_{2}]}=(-1)\otimes\overline{[w^{2}]}=2(-1)\otimes\overline{[w]}=(-1)^{2}\otimes\overline{[w]}=0.$
Thus $\tau$ must be trivial. Hence by Theorem 1.1 we obtain the exact sequence
@@
$H_{4}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\Gamma(\mu_{2}(A))$
is trivial. Since this map is trivial when $\mu_{2}(A)\simeq\mu_{2}(F)=1$, we may assume that $\mu_{2}(A)\simeq\mu_{2}(F)=\{\pm 1\}\simeq\mathbb{Z}/2$. Let $\overline{F}$ be the algebraic closure of $F$. Then $H_{1}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})=0$ and $H_{2}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$ is isomorphic to $K_{2}^{M}(\overline{F})$, the second Milnor $K$-group of $\overline{F}$ [27, page 286, 2.12]. It is well-known that $K_{2}^{M}(\overline{F})$ is uniquely divisible (see [3, Corollary 1.3]) and thus $\mu_{2}(\overline{F})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})=0$. So we have the exact sequence
$\Gamma(\mu_{2}(\overline{F}))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow 0.$
The surjective map $p:\mathit{{\rm SL}}_{2}(\overline{F})\rightarrow\mathit{{\rm PSL}}_{2}(\overline{F})$ induces the commutative diagram with exact rows
@@
The first row is the classical Bloch-Wigner exact sequence [10, App. A], [30, Theorem 5.2]. The second exact sequence can be proved as the first exact sequence, where $\widetilde{\mu}(\overline{F}):=\mu(\overline{F})/\mu_{2}(\overline{F})$ (also see Theorem 4.3 below). The left hand side square is obtained from the commutative diagram
[…]
where for a ring $R$, $T(R):=\bigg{\{}{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}|a\in R^{\times}\bigg{\}}$, $\mathit{{\rm PT}}(R):=T(R)/\mu_{2}(R)I_{2}$ and the middle square is induced by the map $\mu(R)\rightarrow T(R)$, $a\mapsto{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}$. The extension $1\rightarrow\mu_{2}(\overline{F})\rightarrow T(\overline{F})\rightarrow\mathit{{\rm PT}}(\overline{F})\rightarrow 1$ is isomorphic with the extension $1\rightarrow\mu_{2}(\overline{F})\rightarrow\overline{F}^{\times}\overset{()^{2}}{\longrightarrow}\overline{F}^{\times}\rightarrow 1$. The map $\phi=()^{2}:\overline{F}^{\times}\rightarrow\overline{F}^{\times}$ induces the map $\phi_{\ast}:H_{3}(\overline{F}^{\times},\mathbb{Z})\rightarrow H_{3}(\overline{F}^{\times},\mathbb{Z})$. By Corollary 1.4 this leads to multiplication by $2^{2}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))$. This implies that $\ker(\alpha)\simeq\mathbb{Z}/4$ and thus we have the exact sequence
Now the first claim follows from the following commutative diagram with exact rows
[…]
@@
is trivial. Let denote this map by $\phi_{\ast}$. The group $H_{3}(\langle w\rangle,\mathbb{Z})$ is cyclic of order 4 and is generated by
@@
(see [32, Proposition 8.1], [25, Proposition 3.25]), where $B_{\bullet}(\langle w\rangle)\rightarrow\mathbb{Z}$ is the bar resolution of $\mathbb{Z}$ over $\langle w\rangle$. We have
$\phi_{\ast}(\chi)=\overline{[w|w|w]+[w|\!-\!w|w]}=\overline{2[w|w|w]}.$
Since $d_{4}([w|w|w|w])={2[w|w|w]}$ in $B_{\bullet}(\mathit{{\rm PSL}}_{2}(A))_{\mathit{{\rm PSL}}_{2}(A)}$, we have $\phi_{\ast}(\chi)=0$. This shows that the above composite is trivial. Now by the commutative diagram
[…]
@@
Remark 2.3.
is trivial [13, page 103]. The proof of the above theorem shows that if $\mu_{2}(A)=\{\pm 1\}$, then
is a weak stem-extension. But in general it is not a stem-extension (for example take $A=\mathbb{Z}$ or see Proposition 3.1 below).
Corollary 2.4.
is exact.
Proof.
If $|A/\mathfrak{m}_{A}|>3$, then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})=0$ and if $|A/\mathfrak{m}_{A}|=3$, then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq A/\mathfrak{m}_{A}\simeq\mathbb{Z}/3$ [22, Proposition 4.1]. Thus
$\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})=0,\ \ \ \ {{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(A),H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}))=0.$
Hence the claim follows from the above theorem. ∎
Example 2.5.
Let $A$ be a local domain such that $A/\mathfrak{m}_{A}\simeq\mathbb{F}_{2}$. Then $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq A/\mathfrak{m}_{A}^{2}$ [22, Proposition 4.1]. By a careful analysis of the composition
Observe that
If $A=\mathbb{Z}_{(2)}=\{a/b\in\mathbb{Q}:a,b\in\mathbb{Z},2\nmid b\}$ (or $A=\mathbb{Z}_{2}$, the $2$-adic integers), then $2\notin\mathfrak{m}_{A}^{2}$ (by the Nakayama lemma). Thus the sequence
is exact. Note that $H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\simeq\mathbb{Z}/4$ [22, Example 4.2]. With an argument similar to the proof of Proposition 3.1 below, one can show that $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\simeq H_{2}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})$ and the sequence
is exact. We have similar results for $\mathbb{Z}_{2}$. Hence Corollary 2.4 is not necessarily true when $|A/\mathfrak{m}_{A}|=2$.
Example 2.6.
(i) Let $\mathbb{F}_{q}$ be a finite field of odd characteristic. It is well-known that $H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})=0$ when $q\neq 3$ and $H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{3}),\mathbb{Z})\simeq\mathbb{Z}/3$. Moreover, $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})=0$ when $q\neq 9$ and $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{9}),\mathbb{Z})\simeq\mathbb{Z}/3$ [27, page 280]. These facts together with Theorem 2.1 give us the exact sequence
$0\rightarrow\mu_{2}(\mathbb{F}_{q})^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{F}_{q}),\mathbb{Z})\rightarrow 0$
For fields of characteristic $2$ the above exact sequence is trivial. Thus this sequence is exact for any finite field.
(ii) If $F$ is quadratically closed, then $H_{2}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})$ is isomorphic to $K_{2}^{M}(F)$, the second Milnor $K$-group of $F$. It is well-known that for such fields $K_{2}^{M}(F)$ is uniquely 2-divisible (see [3, Corollary 1.3]). Thus by Corollary 2.4 we obtain the exact sequence
This result first proved by Dupont and Sah in [10, A6, App. A].
(iii) Let $F$ be a real closed field. It is mentioned in [11, page 230] that the sequence
is exact and they refer to the proof of the case of quadratically closed fields discussed in [10, App. A]. We couldn’t really verify this. But in the next section we will give a new proof of this fact (see Corollary 4.7). The same result follows from the results of [7].
Consider the product map
$\rho_{\ast}:\mu_{2}(A)\otimes H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}).$
By the above example, this product map is trivial for finite fields, quadratically closed fields and real closed field. In the introduction we asked if this is true for any local domain such that $A/\mathfrak{m}_{A}\neq\mathbb{F}_{2}$. The following example shows that this map is not necessary trivial over more general domains (see also Example 3.2 below).
Example 2.7.
Let $\mathcal{O}_{-3}$ be the ring of algebraic integers of the imaginary quadratic field $\mathbb{Q}(\sqrt{-3})$. The homology groups of $\mathit{{\rm SL}}_{2}(\mathcal{O}_{-3})$ and $\mathit{{\rm PSL}}_{2}(\mathcal{O}_{-3})$ have been calculated in [2, Theorem, page 376] and [28, Theorem 5.7]. We list these homology groups for $1\leq n\leq 3$:
From Theorem 2.1 we have the exact sequence
is non-trivial.
@@
It is known that $\mathit{{\rm SL}}_{2}(\mathbb{Z})\simeq\mathbb{Z}/4\ast_{\mathbb{Z}/2}\mathbb{Z}/6$ and $\mathit{{\rm PSL}}_{2}(\mathbb{Z})\simeq\mathbb{Z}/2\ast\mathbb{Z}/3$. By the Mayer-Vietoris exact sequence associated to amalgamated product [6, Corollary 7.7, Chap. II] one obtains

### v3

@@ 1. The low dimensional homology of central extensions
2. The bloch-wigner exact sequence
Let $A$ be a commutative ring and set $\mathcal{W}_{A}:=\{a\in A:a(1-a)\in{A^{\times}}\}$. The scissors congruence group $\mathcal{P}(A)$ of $A$ is defined as the quotient of the free abelian group generated by symbols $[a]$, $a\in\mathcal{W}_{A}$, by the subgroup generated by the elements
$[a]-[b]+\bigg{[}\frac{b}{a}\bigg{]}-\bigg{[}\frac{1-a^{-1}}{1-b^{-1}}\bigg{]}+\bigg{[}\frac{1-a}{1-b}\bigg{]},$
where $a,b,a/b\in\mathcal{W}_{A}$. Let $S_{\mathbb{Z}}^{2}({A^{\times}}):=({A^{\times}}\otimes_{\mathbb{Z}}{A^{\times}})/\langle a\otimes b+b\otimes a:a,b\in{A^{\times}}\rangle$. The map
$\lambda:\mathcal{P}(A)\rightarrow S_{\mathbb{Z}}^{2}({A^{\times}}),\ \ \ \ [a]\mapsto a\otimes(1-a),$
is well-defined and its kernel is called the Bloch group of $A$ and is denoted by $\mathcal{B}(A)$.
For any local ring $A$ there is a natural map of graded rings $K_{\bullet}^{M}(A)\to K_{\bullet}(A)$, from Milnor $K$-theory to $K$-theory. It is easy to see that
$K_{0}^{M}(A)\simeq K_{0}(A)\simeq\mathbb{Z},\ \ \ \ K_{1}^{M}(A)\simeq K_{1}(A)\simeq{A^{\times}}.$
Moreover, Matsumoto proved that for a field $F$
$K_{2}(F)\simeq K_{2}^{M}(F)\simeq({A^{\times}}\otimes_{\mathbb{Z}}{A^{\times}})/\langle a\otimes(1-a):a\in\mathcal{W}_{F}\rangle$
(see [23, Theorem 4.3.15]).
The indecomposable $K_{3}^{\rm ind}(F)$ of a field $F$ is defined as follows:
$K_{3}^{\rm ind}(F):=K_{3}(F)/K_{3}^{M}(F).$
The Bloch group and the indecomposable part of the third $K$-group are deeply connected.
Theorem 2.1 (Bloch-Wigner exact sequence).
For any field $F$, we have a natural exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))^{\sim}\rightarrow K_{3}^{\rm ind}(F)\rightarrow\mathcal{B}(F)\rightarrow 0$
Proof.
The case of infinite fields has been proved in [27, Theorem 5.2] and the case of finite fields has been settled in [15, Corollary 7.5]. ∎
For any filed $F$, there are natural maps $H_{2}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow K_{2}(F)$ and $H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow K_{3}^{\rm ind}(F)$ (see [15, §5]).
Proposition 2.2.
If $F$ is a quadratically closed field, then we have natural isomorphisms
$K_{2}(F)\simeq K_{2}^{M}(F)\simeq H_{2}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z}),\ \ \ \ \ K_{3}^{\rm ind}(F)\simeq H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z}).$
Furthermore, $K_{2}(F)$ is uniquely $2$-divisible.
Proof.
See [24, page 286, 2.12], [17, Proposition 6.4] and [3, Proposition 1.2]. ∎
The following classical Bloch-Wigner exact sequence will be used in the proof of Theorem 3.1.
Corollary 2.3 (Bloch-Wigner, Dupont-Sah).
If $F$ is a quadritically closed field, then we have the exact sequence
${0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow\mathcal{B}(F)\rightarrow 0.}$
Proof.
For a direct proof see [10, Appedix A]. See also [27, Theorem 5.2]. ∎
Note that the map $\mu(F)\rightarrow\mathit{{\rm SL}}_{2}(F)$ given by $a\mapsto{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}$ induces the natural map
$H_{3}(\mu(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z}).$
By Proposition 1.3, $H_{3}(\mu(F),\mathbb{Z})\simeq{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))$. Now the left map in the above classical Bloch-Wigner exact sequence is the map
${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))\simeq H_{3}(\mu(F),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z}).$
Proposition 2.4 (Coronado [7]).
Let $F$ be a real closed field. Then we have the Bloch-Wigner exact sequence
${0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(F),\mu(F))^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\rightarrow\mathcal{B}(F)\rightarrow 0.}$
In particular, $H_{3}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})\simeq K_{3}^{\rm ind}(F)$
Proof.
See [7, Theorem 6.1] and its proof. ∎
@@ 3. Low dimensional homology of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$
where $\mu_{2}(A)=\{a\in{A^{\times}}:a^{2}=1\}$ and the map $i$ is given by $a\mapsto{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a\end{array}\!\!\right)}={\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}$. Let
$\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}):=\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\rho_{\ast}\Big{(}\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\oplus{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(A),H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}))\Big{)}},$
where $\rho:\mu_{2}(A)\times\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A)$ is the product map and is given by $(a,X)\mapsto aX$. We recall that for a field $F$,
$\mu_{2}(F)^{\sim}\simeq\begin{cases}0&\text{if ${\rm char}(F)=2$}\\ \mathbb{Z}/4&\text{if ${\rm char}(F)\neq 2$}.\end{cases}$
Now we are ready to state and prove our first main result.
@@
Let $A$ be a ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu_{2}(A)\simeq\mu_{2}(F)$. Then we have the $9$-term exact sequence
@@
First we study the homomophism $\tau:\Gamma(\mu_{2}(A))\rightarrow\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ (see (1.1)). Note that by Lemma 1.2, $\Gamma(\mu_{2}(A))\simeq\mu_{2}(A)^{\sim}$. The map $\tau$ is induced by the quadratic map
@@
where $i_{\ast}:\mu_{2}(A)\simeq H_{1}(\mu_{2}(A),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is given by $-1\mapsto{\boldsymbol{c}}(-I_{2})$. Since $w^{2}=-I_{2}$, we have
$(-1)\otimes i_{\ast}(-1)=(-1)\otimes{\boldsymbol{c}}(-I_{2})=(-1)\otimes{\boldsymbol{c}}(w^{2})=(-1)\otimes 2{\boldsymbol{c}}(w)=(-1)^{2}\otimes{\boldsymbol{c}}(w)=0.$
Thus $\tau$ must be trivial. Now by Theorem 1.1 we obtain the exact sequence
@@
$H_{4}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\Gamma(\mu_{2}(A))\simeq\mu_{2}(A)^{\sim}$
is trivial. Since this map is trivial when $\mu_{2}(A)\simeq\mu_{2}(F)=1$, we may assume that $\mu_{2}(A)\simeq\mu_{2}(F)=\{\pm 1\}\simeq\mathbb{Z}/2$.
Let $\overline{F}$ be the algebraic closure of $F$. Then $H_{1}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})=0$ and $H_{2}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$ is uniquely 2-divisible by Proposition 2.2. These imply that ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(\overline{F}),H_{1}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z}))$ and $\mu_{2}(\overline{F})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$ are trivial. Hence by Theorem 1.1, we have the exact sequence
(3.1) $\Gamma(\mu_{2}(\overline{F}))\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})\rightarrow 0.$
Observe that $\Gamma(\mu_{2}(\overline{F}))\simeq\mu_{2}(\overline{F})^{\sim}\simeq\mathbb{Z}/4$ (Lemma 1.2). The surjective map $p:\mathit{{\rm SL}}_{2}(\overline{F})\rightarrow\mathit{{\rm PSL}}_{2}(\overline{F})$ induces the commutative diagram with exact rows
00{0}${{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))}$${H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})}$${\mathcal{\mathcal{B}}(\overline{F})}$00{0}00{0}${{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(\overline{F}),\widetilde{\mu}(\overline{F}))}$${H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})}$${\mathcal{B}(\overline{F})}$${0.}$$\scriptstyle{\alpha}$$\scriptstyle{p_{\ast}}$
The first row is the classical Bloch-Wigner exact sequence discussed in the previous section (Corollary 2.3). The second exact sequence can be proved as the first exact sequence, where $\widetilde{\mu}(\overline{F}):=\mu(\overline{F})/\mu_{2}(\overline{F})$ (see [10, Theorem 4.10 and Appendix A]. Also see [20, Corollary 3.5] below). The left hand side square is obtained from the commutative diagram
${{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))}$${H_{3}(\mu(\overline{F}),\mathbb{Z})}$${H_{3}({\rm T}(\overline{F}),\mathbb{Z})}$${H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})}$${{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(\overline{F}),\widetilde{\mu}(\overline{F}))}$${H_{3}(\widetilde{\mu}(\overline{F}),\mathbb{Z})}$${H_{3}(\mathit{{\rm PT}}(\overline{F}),\mathbb{Z})}$${H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z}),}$$\scriptstyle{\alpha}$$\scriptstyle{p_{\ast}}$$\scriptstyle{p_{\ast}}$$\scriptstyle{p_{\ast}}$
where for a ring $R$,
${\rm T}(R):=\bigg{\{}{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}|a\in R^{\times}\bigg{\}},\ \ \ \ \ \mathit{{\rm PT}}(R):={\rm T}(R)/\mu_{2}(R)I_{2}$
and the middle square is induced by the map $\mu(R)\rightarrow{\rm T}(R)$, $a\mapsto{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}$. The extension
$1\rightarrow\mu_{2}(\overline{F})\rightarrow{\rm T}(\overline{F})\rightarrow\mathit{{\rm PT}}(\overline{F})\rightarrow 1$
is isomorphic with the extension
$1\rightarrow\mu_{2}(\overline{F})\rightarrow\overline{F}^{\times}\overset{()^{2}}{\longrightarrow}\overline{F}^{\times}\rightarrow 1$
The map $\phi=()^{2}:\overline{F}^{\times}\rightarrow\overline{F}^{\times}$ induces the map $\phi_{\ast}:H_{3}(\overline{F}^{\times},\mathbb{Z})\rightarrow H_{3}(\overline{F}^{\times},\mathbb{Z})$. By Proposition 1.3 we have the split exact sequence
$\begin{array}[]{c}0\rightarrow\bigwedge_{\mathbb{Z}}^{3}\overline{F}^{\times}\rightarrow H_{3}(\overline{F}^{\times},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))\rightarrow 0.\end{array}$
By Corollary 1.4 this leads to multiplication by $2^{2}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))$. Since
${{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F}))\simeq\mu(\overline{F}),$
the above fact implies that $\ker(\alpha)\simeq\mathbb{Z}/4$. On the other hand by the Snake lemma $\ker(\alpha)\simeq\ker(p_{\ast})$. Thus the left map of the exact sequence (3.1) must be injective and so we have the exact sequence
Observe that
$\Gamma(\mu_{2}(\overline{F}))\simeq\mu_{2}(\overline{F})^{\sim}\hookrightarrow\mu(\overline{F})^{\sim}\simeq\mu(\overline{F})\simeq{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(\overline{F}),\mu(\overline{F})).$
The composite $A\rightarrow F\hookrightarrow\overline{F}$ induces the morphism of extensions
@@
Now by Theorem 1.1 we obtain the commutative diagram with exact rows
${H_{4}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})}$${\Gamma(\mu_{2}(A))}$${\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}$${H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})}$00{0}${\Gamma(\mu_{2}(\overline{F}))}$${H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})}$${H_{3}(\mathit{{\rm PSL}}_{2}(\overline{F}),\mathbb{Z})}$${0,}$
where the exactness of the second row has just been shown. Since $\mu_{2}(A)\simeq\mu_{2}(F)=\mu_{2}(\overline{F})$, we have
$\Gamma(\mu_{2}(A))\simeq\Gamma(\mu_{2}(\overline{F})).$
Now from the above diagram it follows that the map $\Gamma(\mu_{2}(A))\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is injective. This complete the proof of the first claim.
@@
is trivial. Let denote this map by $\varphi_{\ast}$. The group $H_{3}(\langle w\rangle,\mathbb{Z})$ is cyclic of order 4 and is generated by
@@
(see [29, Proposition 8.1] or [22, Proposition 3.25]), where $B_{\bullet}(\langle w\rangle)\rightarrow\mathbb{Z}$ is the bar resolution of $\mathbb{Z}$ over $\langle w\rangle$. We have
$\varphi_{\ast}(\chi)=\overline{[w|w|w]+[w|\!-\!w|w]}=\overline{2[w|w|w]}.$
Since
$d_{4}([w|w|w|w])={2[w|w|w]}$
in $B_{\bullet}(\mathit{{\rm PSL}}_{2}(A))_{\mathit{{\rm PSL}}_{2}(A)}$, we have $\varphi_{\ast}(\chi)=0$. This shows that the above composite is trivial. Now by the commutative diagram
${H_{3}(\langle w\rangle,\mathbb{Z})}$${\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}$${H_{3}(\langle w\rangle,\mathbb{Z})}$${H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z}),}$
it is sufficient to prove that the map $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$ is injective. Let $\mathfrak{i}^{2}=-1$. Then $w$ is conjugate to ${\left(\!\!\begin{array}[]{cc}\!\!\mathfrak{i}&\!\!0\\ \!\!0&\!\!\mathfrak{i}^{-1}\end{array}\!\!\right)}$. It is known by the classical Bloch-Wigner exact sequence (Corollary 2.3) that the map
$H_{3}(\mu(\overline{F}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$
which is induced by $\mu(\overline{F})\rightarrow\mathit{{\rm SL}}_{2}(\overline{F})$, $a\mapsto{\left(\!\!\begin{array}[]{cc}\!\!a&\!\!0\\ \!\!0&\!\!a^{-1}\end{array}\!\!\right)}$, is injective. Thus the map
$H_{3}(\langle\mathfrak{i}\rangle,\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\overline{F}),\mathbb{Z})$
is injective. This proves our claim. ∎
Remark 3.2.
The above theorem is also valid if we replace $\mathit{{\rm SL}}_{2}(A)$ with the elementary subgroup ${\rm E}_{2}(A)$. The proof stays the same. Just one should note that over a field $F$, ${\rm E}_{2}(F)=\mathit{{\rm SL}}_{2}(F)$.
@@
It is known that $\mathit{{\rm SL}}_{2}(\mathbb{Z})\simeq\mathbb{Z}/4\ast_{\mathbb{Z}/2}\mathbb{Z}/6$ and $\mathit{{\rm PSL}}_{2}(\mathbb{Z})\simeq\mathbb{Z}/2\ast\mathbb{Z}/3$. By the Mayer-Vietoris exact sequence associated to a amalgamated product [6, Corollary 7.7, Chap. II] one obtains